# TAD-Bench: A Comprehensive Benchmark for Embedding-Based Text Anomaly Detection

### v1

@@ 4.3 Anomaly Detectors
The embeddings derived from these models were subsequently used as input features for anomaly detection algorithms. To identify anomalous instances, we employed 8 state-of-the-art anomaly detection methods sourced from the PyOD library [^1] PyOD: https://pyod.readthedocs.io/en/latest/index.html (Zhao et al., 2019). These algorithms include KNN, LOF, OCSVM, iForest, INNE, ECOD, HBOS and COPOD . These algorithms were selected to capture diverse anomaly detection paradigms, ensuring robust detection across varying dataset characteristics.
For reproducibility and consistency, the default hyperparameter settings provided in the respective algorithm implementations and original papers were adopted. This approach minimizes user bias and allows for a fair comparison of algorithm performance when applied to the embeddings generated by the different models. The combination of diverse embedding models and anomaly detection algorithms ensures a comprehensive evaluation of text anomaly detection in terms of both computational efficiency and detection effectiveness.
Models Max Tokens # Dimensions # Parameters
BERT 512 768 110 M
MINILM 512 384 22.7 M
O-ada 8191 1536 350 M
O-small 8191 1536 125 M
O-large 8191 3072 1.3 B
LLAMA 4096 2048 1 B
stella 2048 1024 400 M
Qwen 8192 1536 1.5 B
@@ 4.4 Evaluation Criteria and Trials
Performance was evaluated using the Area Under the Receiver Operating Characteristic Curve (AUROC), which is a widely used criteria in anomaly detection tasks. Each experiment was conducted 5 times, and the average AUROC was reported to ensure result reliability.
The entire pipeline, including embedding extraction and anomaly detection, was implemented in PyTorch. Experiments were executed on a computational setup equipped with a Ryzen 9 5900X 12-core CPU for data preprocessing and model orchestration, and an Nvidia RTX 3060 GPU with 12GB of memory for model inference and embedding generation.
@@ 5 Experiments
Table 3: Evaluation across 6 datasets in terms of ROC-AUC.
@@
Figure 1: Boxplot of AUCROC scores for anomaly detectors on different embeddings across six datasets.
5.1 Applicability of LLM Embeddings with Anomaly Detectors
Table 1 summarizes the performance of various anomaly detectors combined with LLM-derived embeddings across different datasets, while Figure 1 highlights their strong performance in specific tasks, particularly spam detection. In both the email spam and SMS spam detection tasks (Figure 1(a) and Figure 1(d), many embedding-detector combinations achieve high AUC scores, with several exceeding 0.8. This strong performance can be attributed to the explicit nature of spam-related features, such as the presence of URLs, nonsensical text, or repetitive patterns. An example from the Email Spam dataset is shown below:
Subject: oxyccontttin no script needeeed your place to ggo too for all ur prreexxxxiscrlpt 10 n pi sx , paaaaain killerzxss noeoo presscippt http : / / hyyydroccodeeeine vicccodinne / vic geeet reeeliefff noowee http : / / offfmeebabyy
These features are effectively represented in the semantic spaces created by general-purpose embeddings, enabling anomaly detectors to distinguish spam messages from legitimate ones. Additionally, the relatively small variance in detection performance across embeddings suggests that spam detection relies on surface-level linguistic patterns that are well captured by the embeddings used in this study.
In contrast, the performance on hate speech and offensive language detection tasks (Figure 1(c) and Figure 1(f)) is consistently weaker, with AUC scores rarely exceeding 0.6 across embedding-detector combinations. This suggests that the embeddings struggle to capture the nuanced and context-dependent features necessary for these tasks. For example, hate speech often involves implicit cues such as sarcasm, cultural references, or indirect hostility, which may not be adequately represented in the embeddings. Similarly, offensive language detection, as seen in the OLID dataset, requires distinguishing fine-grained differences in tone or intent, such as the distinction between neutral, offensive, or subjective language. These distinctions often depend on broader contextual information, such as the discourse or dialogue in which the language appears. These distinctions often depend on broader contextual information, such as the discourse or dialogue in which the language appears. For example, without additional context, such as the speaker’s intent or the conversational background, the following statement from OLID dataset remains ambiguous whether this statement qualifies as hate speech:
@USER #metoo are all racist!
@@ 5.2 Comparative Effectiveness of Embeddings in Anomaly Detection
In comparison, other embeddings, such as MINILM, exhibit strong performance in specific tasks but lack consistency across more complex datasets. For instance, MINILM achieves exceptional AUC scores of 0.9526 and 0.9626 on the Email Spam datasets, respectively, when used with INNE and OCSVM. However, its performance drops significantly on datasets like OLID and LIAR2, indicating limitations in handling tasks where deeper contextual or domain-specific cues are required. Similarly, embeddings such as stella and Qwen show moderate performance, excelling in a limited subset of tasks but failing to achieve the same level of versatility as the OpenAI embeddings.
These observations suggest that the OpenAI embeddings, particularly O-small and O-large, provide the most robust and consistent performance across a wide range of tasks. Their ability to effectively represent both explicit textual features (e.g., spam detection tasks) and subtle contextual variations (e.g., Covid Fake News and OLID) underscores their suitability for anomaly detection scenarios that require both surface-level semantic extraction and deeper linguistic understanding.
Moreover, the collective performance of the OpenAI family emphasizes the adaptability and efficacy of transformer-based embeddings in addressing anomaly detection challenges across datasets with diverse linguistic and semantic complexities. OpenAI embeddings consistently emerge as dominant and versatile solutions in various scenarios, making them highly effective for a wide range of applications.
@@ 5.3 Performance Across Anomaly Detectors
Figure 2: Average rank (lower the better) of 3 differernt OpenAI embeddings-based methods on AUCROC across 6 datasets
@@
Overall, kNN, INNE, and ECOD demonstrate robust and consistent performance across embeddings and tasks, making them well-suited for general-purpose anomaly detection pipelines. In contrast, methods like LOF, COPOD, and iForest show weaker adaptability, likely due to their limitations in handling high-dimensional and semantically rich embeddings.
@@ 6 Conclusion
In this work, we present a comprehensive benchmark for embedding-based anomaly detection in natural language processing, evaluating the interplay between large language model (LLM) embeddings and classical anomaly detection algorithms across three diverse domains: spam detection, fake news detection, and offensive language detection. Our results reveal both the strengths and limitations of LLM embeddings, demonstrating their effectiveness in tasks with explicit patterns but highlighting challenges in capturing implicit, context-dependent anomalies.
Limitations and Future Works
@@
In addition, our work uses pre-trained LLM embeddings and default hyperparameters for anomaly detectors, ensuring consistency but potentially underestimating their best-case performance. Fine-tuning LLMs on domain-specific data could improve embedding quality, while systematic hyperparameter optimization might unlock the full potential of anomaly detectors. Future research should explore these directions, leveraging techniques like AutoML to streamline both embedding fine-tuning and parameter tuning, thereby achieving more competitive performance.
Moreover, TAD-Bench focuses solely on embedding-based methods, excluding end-to-end approaches that directly process raw text. While embeddings offer modularity and efficiency, end-to-end models like autoencoders or transformer-based systems may capture richer contextual information and handle more complex anomalies. Future work should incorporate end-to-end models and explore hybrid approaches that combine the strengths of both paradigms, providing a more comprehensive evaluation of anomaly detection methods in NLP.
@@
All the data and information in this paper are from public source.

### v2

@@ 3.3 Anomaly Detectors
The embeddings derived from these models were subsequently used as input features for anomaly detection algorithms. To identify anomalous instances, we employed 8 anomaly detection methods sourced from the PyOD library [^1] PyOD: https://pyod.readthedocs.io/en/latest/index.html Zhao et al. (2019). These algorithms include KNN, LOF, OCSVM, iForest, INNE, ECOD, HBOS and COPOD . These algorithms were selected to capture diverse anomaly detection paradigms, ensuring robust detection across datasets with varying characteristics, structures, and distributions.
For reproducibility and consistency, we implemented all algorithms using their default hyperparameter settings as specified in their original implementations and research papers. This approach minimizes subjective bias and enables fair comparison across different embedding models. Additionally, we conducted comprehensive grid search for optimal hyperparameter configurations, with search ranges detailed in Table 4. By evaluating both default and optimized settings across our diverse collection of embedding models and detection algorithms, we provide a thorough assessment of text anomaly detection performance in terms of both computational efficiency and detection effectiveness.
@@ 3.4 Evaluation Criteria and Trials
Performance was evaluated using the Area Under the Receiver Operating Characteristic Curve (AUROC), a widely adopted metric in anomaly detection tasks for measuring the trade-off between true positive and false positive rates. To ensure the reliability and robustness of the results, each experiment was repeated 5 times, and the average AUROC score was reported.
@@ 4 Experiments
4.1 Domain Generalization
Table 5 summarizes the performance of various anomaly detectors combined with LLM-derived embeddings across different datasets, while Figure 1 highlights their strong performance in specific tasks, particularly spam detection. In both the email spam and SMS spam detection tasks (Figure 5(a) and Figure 5(d), many embedding-detector combinations achieve high AUC scores, with several exceeding 0.8. This strong performance can be attributed to the explicit nature of spam-related features, such as the presence of URLs, nonsensical text, or repetitive patterns. An example Case 1 from the Email Spam dataset is shown below:
Case 1: Subject: oxyccontttin no script needeeed your place to ggo too for all ur prreexxxxiscrlpt 10 n pi sx , paaaaain killerzxss noeoo presscippt http : / / hyyydroccodeeeine vicccodinne / vic geeet reeeliefff noowee http : / / offfmeebabyy
Figure 3: t-SNE demonstration of Case 1 (green star) embedding extracted by O-large.
Figure 3 shows that case 1 is located at the edge of the embedding space and deviates from the main data distribution, thus making it easy to be detected.
These features are effectively represented in the semantic spaces created by general-purpose embeddings, enabling anomaly detectors to distinguish spam messages from legitimate ones. Additionally, the relatively small variance in detection performance across embeddings suggests that spam detection primarily relies on surface-level linguistic patterns, which are effectively captured by the embeddings employed in this study.
In contrast, the performance on hate speech and offensive language detection tasks (Figure 5(c) and Figure 5(f)) is consistently weaker, with AUC scores rarely exceeding 0.6 across embedding-detector combinations. This suggests that the embeddings struggle to capture the nuanced and context-dependent features necessary for these tasks. For instance, hate speech often relies on implicit cues such as sarcasm, cultural references, or subtle forms of hostility, which may not be fully captured by standard embeddings. Similarly, offensive language detection, as observed in the OLID dataset, requires identifying fine-grained differences in tone, intent, and subjectivity, such as distinguishing between neutral, offensive, and sarcastic expressions. These distinctions often depend on broader contextual information, such as the discourse or dialogue in which the language appears.
For example, without additional context, such as the speaker’s intent or the conversational background, the following statement from OLID dataset remains ambiguous whether this statement qualifies as hate speech:
Case 2: @USER #metoo are all racist!
Figure 4: t-SNE demonstration of Case 2 (green star) embedding extracted by O-large.
Figure 4 shows that case 2 is mixed in the normal data distribution, making it difficult to be detected.
Table 2: Evaluation across 6 datasets in terms of AUROC.
@@ 4.1 Domain Generalization
Figure 5: Boxplot of AUCROC scores for anomaly detectors on different embeddings across 6 datasets.
@@ 4.2 Comparative Effectiveness of Embeddings in Anomaly Detection
In comparison, other embeddings, such as MINILM, exhibit strong performance in specific tasks but lack consistency across more complex datasets. For instance, MINILM achieves exceptional AUC scores of 0.9526 and 0.9626 on the Email Spam datasets when paired with INNE and OCSVM, respectively. However, its performance declines significantly on datasets like OLID and LIAR2, suggesting limitations in capturing deeper contextual or domain-specific cues essential for these tasks. Similarly, embeddings such as stella and Qwen exhibit moderate performance, excelling in a limited subset of tasks but failing to match the versatility of OpenAI embeddings. Their inconsistent performance across datasets indicates that while they may effectively capture certain linguistic patterns, they struggle with tasks requiring a broader understanding of context, intent, or nuanced semantics.
These observations suggest that OpenAI embeddings, deliver the most robust and consistent performance across a diverse set of tasks. Their ability to effectively capture both explicit textual features (e.g., in spam detection) and nuanced contextual variations (e.g., in Covid Fake News and OLID) highlights their versatility. This underscores their suitability for anomaly detection scenarios that demand both surface-level pattern recognition and deeper linguistic comprehension, making them well-equipped for handling a wide range of text-based anomalies.
@@ 4.3 Performance Across Anomaly Detectors
Figure 6: Average rank (lower the better) of 3 differernt OpenAI embeddings-based methods on AUCROC across 6 datasets.
@@ 5 Conclusion
In this paper, we present a comprehensive benchmark for embedding-based text anomaly detection, systematically evaluating the interplay between LLM embeddings and classical anomaly detection algorithms across three diverse domains. Our results reveal both the strengths and limitations of embedding-based anomaly detection methods, demonstrating their effectiveness in tasks with explicit and well-defined patterns while highlighting challenges in capturing implicit, context-dependent anomalies that require broader contextual cues. These findings emphasize the need for more adaptive embeddings and hybrid detection strategies that integrate external knowledge and contextual reasoning.
Limitations
@@
Moreover, TAD-Bench focuses solely on embedding-based methods, excluding end-to-end approaches that directly process raw text because due to modularity and efficiency of embedding-based methods, and NLP-ADBench has also shown better performance of embedding-based methods than end-to-end methods. However, end-to-end models like autoencoders or transformer-based methods may capture richer contextual information and handle more complex anomalies. Future work should incorporate end-to-end models and explore hybrid approaches that combine the strengths of both paradigms, providing a more comprehensive evaluation of anomaly detection methods in NLP.
@@
This study adheres to ethical research practices and considerations in the development and evaluation of text anomaly detection methods.
Use of Potentially Offensive Language. Some examples in this paper may contain offensive, harmful, or misleading language. These examples are used purely for illustrative purposes to demonstrate the challenges of text anomaly detection in real-world scenarios. They do not reflect the opinions, beliefs, or endorsements of the authors.
Data Sources and Usage. All datasets used in this study are sourced from publicly available research datasets that have been previously used in NLP and anomaly detection research. Proper citations and references to the original datasets are provided in the paper. No private, proprietary, or personally identifiable information was used in this study.
Risks and Responsible Use. Because anomaly detection models can be misused for purposes such as censorship, surveillance, or unfair content moderation. We strongly emphasize that our benchmark is intended for research and academic purposes only and should be used responsibly with consideration of ethical and societal implications.
Use of AI Assistance We acknowledge the use of AI-based writing assistants for grammar refinement, spelling correction, and improving the clarity of our manuscript. However, all intellectual contributions, experimental designs, analyses, and conclusions in this paper are solely the work of the authors.
Appendix A Problem Definitions
Let $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$ represent a corpus consisting of $N$ textual instances, where each instance $x_{i}\in\mathcal{X}$ is represented as a sequence of tokens:
$x_{i}=[t_{1},t_{2},\dots,t_{L_{i}}],$
where $L_{i}$ denotes the sequence length of $x_{i}$. The goal of text anomaly detection is to identify a subset of instances $\mathcal{D}_{\text{anomaly}}\subset\mathcal{D}$, such that $\mathcal{D}_{\text{anomaly}}$ contains samples that deviate significantly from the majority of the dataset $\mathcal{D}_{\text{normal}}=\mathcal{D}\setminus\mathcal{D}_{\text{anomaly}}$.
To achieve this, an anomaly detection algorithm $g$ is applied to the representations of the textual instances to identify potential anomalies. (1) Each text instance $x_{i}$ is first mapped to a fixed-dimensional vector $\mathbf{z}_{i}\in\mathbb{R}^{d}$ using an embedding model $\phi:\mathcal{X}\to\mathbb{R}^{d}$, such that $\mathbf{z}_{i}=\phi(x_{i})$. (2) The anomaly detection algorithm then assigns an anomaly score $s_{i}=g(\mathbf{z}_{i})$ to each instance, $s_{i}\in[0,1]$. Based on a predefined threshold $\tau$, an instance $x_{i}$ is classified as anomalous if:
$x_{i}\in\mathcal{D}_{\text{anomaly}}\iff s_{i}\geq\tau.$
The objective of text anomaly detection is to ensure that $g$ effectively distinguishes between normal and anomalous instances, even in the absence of labeled data, while being robust to the inherent variability and high dimensionality of textual data.
Appendix B Clarification Between Anomaly and Novelty Detection
Text Anomaly Detection (TAD), as defined in Section A, focuses on identifying instances that deviate significantly from the majority of a dataset, regardless of whether anomalies are present during training. While some prior studies (e.g., AD-NLP Bejan et al. (2023), NLP-ADBench Li et al. (2024) and AD-LLM Yang et al. (2024b)) assume training data contains only normal instances and testing data includes both normal and anomalous samples, this setup aligns more closely with novelty detection Pimentel et al. (2014). Novelty detection specifically targets never-before-seen anomalies that are absent from the training phase, often treating anomalies as entirely novel classes.
In contrast, our benchmark evaluates a broader spectrum of anomaly detection scenarios. We do not restrict the training data to purely normal instances, allowing for potential partial supervision or contaminated training sets (e.g., realistic scenarios where anomalies may unintentionally exist in training data). This setup reflects real-world applications where anomaly types are not always fully known a prior, and detection systems must generalize across domains and anomaly types.
This distinction underscores our goal of advancing generalizable anomaly detection systems for real-world NLP applications, where anomalies may exhibit both explicit and context-dependent patterns.
Appendix C Datasets
Table 3: Embedding time of 6 datasets in seconds.
Embeddings Email-Spam SMS-Spam COVID-Fake LIAR2 Hate-Speech OLID
BERT 64.95 s 10.87 s 4.48 s 3.7 s 9.79 s 1.81 s
MINILM 3.58 s 1.48 s 0.64 s 0.52 s 1.50 s 0.45 s
O-ada 154.08 s 33.76 s 17.35 s 16.67 s 35.57 s 8.9 s
O-small 166.73 s 34.16 18.32 16.10 34.09 s 9.0 s
O-large 206.07 s 41.78 s 20.58 s 29.97 s 44.20 s 10.72 s
Llama 545.51 s 129.16 s 38.15 s 28.93 s 71.49 s 18.02 s
stella 99.75 s 19.38 s 12.41 s 9.95 s 20.18 s 5.72 s
Qwen 745.85 s 129.16 s 58.19 s 40.29 s 183.24 s 20.49 s
Email-Spam  [^2] https://huggingface.co/datasets/kendx/NLP-ADBench/tree/main/datasets/email_spam Metsis et al. (2006) contains 5,171 emails labeled as spam or ham, with spam treated as the anomaly class. We utilized the preprocessed version provided in Li et al. (2024).
SMS-Spam  [^3] https://archive.ics.uci.edu/dataset/228/sms+spam+collection Almeida et al. (2011) comprises 5,574 SMS messages originally labeled as spam or ham. Spam messages are designated as the anomaly.
COVID-Fake  [^4] https://github.com/diptamath/covid_fake_news?tab=readme-ov-file Das et al. (2021) comprises posts collected from social media platforms and fact-checking websites. Real news items were sourced from verified outlets providing accurate COVID-19 information, while fake news was gathered from tweets, posts, and articles containing misinformation about COVID-19. Fake news is treated as the anomaly class.
LIAR2  [^5] https://github.com/chengxuphd/liar2?tab=readme-ov-file Xu and Kechadi (2024) consists of approximately 23,000 statements manually labeled by professional fact-checkers for fake news detection tasks. The "True" class, representing accurate statements, is considered the normal class, while the "Pants on Fire" class, representing highly misleading statements, is treated as the anomaly.
OLID  [^6] https://sites.google.com/site/offensevalsharedtask/olid Zampieri et al. (2019b) Zampieri et al. (2019a) contains 14,200 annotated English tweets, categorized using a three-level annotation model. For this benchmark, only the Level A (Offensive Language Detection) annotations are used, where tweets labeled as offensive are considered as anomalies, and non-offensive tweets are considered as normal.
Hate-Speech  [^7] https://github.com/t-davidson/hate-speech-and-offensive-language Davidson et al. (2017) contains tweets annotated by CrowdFlower users. The tweet content is used as data, with "hate speech" treated as anomalies.
Appendix D Evaluation with Hyperparameter Search
Experiments with hyperparameter search reveal significant insights into the performance dynamics of anomaly detection algorithms when paired with various embedding models. The hyperparameters for LOF and HBOS are searched in $\{5,10,20,40\}$. For iForest and iNNE, sample points are searched in $\{2,4,8,16,32,64,128,256\}$ and with defalt $t=200$. $\gamma$ of OCSVM are searched in $\{0.05,0.1,0.5,1\}$ and with RBF kernel. Both COPOD and ECOD don’t have hyperparameters.
When comparing Table 4 with the default parameters results in Table 2, we observe that kNN emerges as a particularly strong performer across multiple embedding models, especially with the OpenAI family of embeddings. This suggests that distance-based approaches effectively leverage the semantic information captured by these models, particularly in tasks like spam and fake news detection.
INNE demonstrates the most balanced and robust performance profile when considering average scores across all datasets and embedding combinations. Its isolation-based approach with hypersphere partitioning appears particularly well-suited to the complex topological structure of embedding spaces, allowing it to identify local anomalies that other methods might miss. The performance improvement of INNE after hyperparameter optimization is especially notable with embedding models like Llama and stella, suggesting a strong complementarity between isolation-based algorithms and these embedding architectures.
Table 4: Evaluation across 6 datasets in terms of AU-ROC under parameter search.
BERT kNN 0.7707 0.5663 0.8467 0.6594 0.5033 0.5164 0.6438
OCSVM 0.5862 0.5322 0.4868 0.5059 0.5154 0.4982 0.5208
IForest 0.7187 0.6307 0.7794 0.6178 0.4957 0.5315 0.6290
LOF 0.7014 0.3531 0.8707 0.676 0.4636 0.5018 0.5944
ECOD 0.7309 0.6235 0.7722 0.6175 0.4889 0.4933 0.6211
INNE 0.7748 0.6551 0.8381 0.6572 0.4897 0.5078 0.6538
HBOS 0.717 0.6252 0.7711 0.6209 0.4935 0.5065 0.6224
COPOD 0.6454 0.5929 0.7714 0.6242 0.4971 0.5189 0.6083
MINILM kNN 0.9669 0.3537 0.8413 0.7249 0.5804 0.5063 0.6623
OCSVM 0.6929 0.582 0.5241 0.5549 0.5538 0.5988 0.5844
IForest 0.9112 0.5655 0.7373 0.5982 0.4582 0.4872 0.6263
LOF 0.6896 0.5507 0.8095 0.6759 0.5286 0.5689 0.6372
ECOD 0.9525 0.5934 0.7581 0.6532 0.3786 0.4208 0.6261
INNE 0.9603 0.6531 0.8226 0.6754 0.5019 0.5381 0.6919
HBOS 0.9489 0.6177 0.7447 0.6614 0.3937 0.4333 0.6333
COPOD 0.9453 0.6317 0.7416 0.6695 0.371 0.4037 0.6271
O-ada kNN 0.9506 0.3535 0.9094 0.7921 0.6341 0.5243 0.6940
OCSVM 0.7535 0.5633 0.56 0.4513 0.5049 0.6474 0.5801
IForest 0.8826 0.7168 0.7493 0.6268 0.5137 0.5194 0.6681
LOF 0.4203 0.536 0.798 0.7568 0.4683 0.5368 0.5860
ECOD 0.938 0.8822 0.815 0.72 0.461 0.4986 0.7191
INNE 0.9469 0.8905 0.8633 0.7558 0.5236 0.5319 0.7520
HBOS 0.9438 0.8851 0.8194 0.7212 0.4692 0.5124 0.7252
COPOD 0.9502 0.8759 0.8153 0.7201 0.4513 0.4811 0.7157
O-small kNN 0.9637 0.229 0.94 0.7872 0.6416 0.5587 0.6867
OCSVM 0.5716 0.5253 0.5165 0.4664 0.5505 0.6059 0.5394
IForest 0.9021 0.6007 0.7871 0.6086 0.4805 0.5388 0.6530
LOF 0.4642 0.5215 0.8596 0.7491 0.4509 0.5602 0.6009
ECOD 0.9481 0.6301 0.8808 0.7022 0.4249 0.5295 0.6859
INNE 0.9572 0.6471 0.9295 0.7374 0.5234 0.5884 0.7305
HBOS 0.953 0.6284 0.8738 0.7009 0.4245 0.5301 0.6851
COPOD 0.9605 0.5722 0.8664 0.6974 0.4017 0.4963 0.6658
O-large kNN 0.9121 0.1698 0.9537 0.7687 0.6291 0.5497 0.6639
OCSVM 0.7794 0.5353 0.5124 0.4302 0.5086 0.5938 0.5600
IForest 0.9142 0.5603 0.8089 0.5887 0.507 0.554 0.6555
LOF 0.5009 0.4748 0.8736 0.7435 0.4296 0.5429 0.5942
ECOD 0.9487 0.6422 0.8875 0.654 0.3959 0.4967 0.6708
INNE 0.9537 0.6765 0.9295 0.7125 0.4899 0.519 0.7135
HBOS 0.9552 0.6562 0.8866 0.6458 0.3882 0.5108 0.6738
COPOD 0.9639 0.6798 0.8854 0.6536 0.3537 0.498 0.6724
Llama kNN 0.9323 0.5389 0.8668 0.7241 0.4991 0.4081 0.6616
OCSVM 0.8929 0.5431 0.5785 0.5045 0.5224 0.529 0.5951
IForest 0.8971 0.7509 0.7815 0.6833 0.4691 0.4267 0.6681
LOF 0.7188 0.461 0.8704 0.7286 0.4564 0.5093 0.6241
ECOD 0.8844 0.7573 0.7819 0.6989 0.4643 0.3998 0.6644
INNE 0.918 0.8083 0.8666 0.7074 0.4778 0.4929 0.7118
HBOS 0.9029 0.7908 0.7782 0.7064 0.4606 0.3961 0.6725
COPOD 0.9153 0.8163 0.7584 0.7291 0.4435 0.3526 0.6692
stella kNN 0.935 0.3451 0.9034 0.6884 0.4746 0.5016 0.6414
OCSVM 0.8655 0.5818 0.5701 0.5071 0.5516 0.5084 0.5974
IForest 0.8884 0.7398 0.7927 0.517 0.4284 0.4716 0.6397
LOF 0.4722 0.5159 0.7749 0.6575 0.4341 0.5598 0.5691
ECOD 0.9075 0.7894 0.8115 0.5023 0.3421 0.4395 0.6321
INNE 0.927 0.8139 0.8466 0.6266 0.4118 0.5243 0.6917
HBOS 0.9178 0.8038 0.8161 0.4952 0.3391 0.4276 0.6333
COPOD 0.93 0.8589 0.8167 0.4936 0.3018 0.3797 0.6301
Qwen kNN 0.9171 0.2671 0.8438 0.6626 0.4991 0.4602 0.6083
OCSVM 0.8304 0.5431 0.5279 0.5105 0.5224 0.5712 0.5843
IForest 0.8758 0.6552 0.7598 0.6169 0.4691 0.4948 0.6453
LOF 0.6915 0.4299 0.8575 0.6972 0.4564 0.4894 0.6037
ECOD 0.8678 0.6648 0.768 0.6172 0.4643 0.4773 0.6432
INNE 0.8989 0.7057 0.8639 0.6733 0.4778 0.518 0.6896
HBOS 0.8877 0.6886 0.7644 0.6187 0.4606 0.4741 0.6490
COPOD 0.9044 0.7393 0.7463 0.6291 0.4435 0.4336 0.6494
Appendix E Embedding Models
Models Max Tokens # Dimensions # Parameters
BERT 512 768 110 M
MINILM 512 384 22.7 M
O-ada 8191 1536 -
O-small 8191 1536 -
O-large 8191 3072 -
LLAMA 4096 2048 1.24 B
stella 2048 1024 435 M
Qwen 8192 1536 1.54 B
To effectively represent textual data, we use various pre-trained embedding models that transform text into dense vector representations. Table 5 summarizes the embedding models employed in this paper. These embeddings serve as feature inputs for anomaly detection models, enabling them to capture semantic similarities and deviations in text. We selected a diverse set of embedding models, balancing between model size, token length limits, and computational efficiency. The models used in this study are:
Beyond model size and token limits, computational efficiency is a key factor in selecting embedding models, particularly for real-world applications where inference speed is critical. Table 3 presents the embedding time (in seconds) required to process six datasets using each embedding model.
From the Table 3, we observe a significant variation in embedding extraction time. MINILM is the fastest across all datasets, taking only a few seconds, making it ideal for applications requiring real-time embedding generation. BERT offers a moderate trade-off, with embedding times significantly lower than larger models but higher than MINILM. OpenAI’s embeddings (O-ada, O-small, O-large) are relatively slow, likely due to their high-dimensional output and extended token support. Llama and Qwen models require the most computation, with Qwen taking up to 745.85 seconds on the Email-Spam dataset, reflecting the high computational cost of large autoregressive models.
Appendix F Comparative Analysis of Anomaly Detection Algorithms
Anomaly detection algorithms vary in their underlying assumptions, computational efficiency, and effectiveness across different types of data distributions. In this section, we provide a comparative analysis of the eight anomaly detection methods used in this study: kNN, OCSVM, iForest, LOF, HBOS, ECOD, INNE and COPOD.
Distance-based methods, such as kNN, define anomalies based on their relative distance to surrounding points. kNN anomaly detection computes the distance between a data point and its $k$th nearest neighbor, with larger distances indicating potential anomalies. This method is conceptually simple and effective in low-dimensional spaces with clear separation between normal and anomalous points. However, its primary drawback is the curse of dimensionality, where distance metrics lose discriminative power as dimensionality increases. Additionally, kNN is computationally expensive, with a worst-case complexity of $O(n^{2})$, making it impractical for large datasets without optimizations such as approximate nearest neighbor search.
Density-based approaches assume that anomalies reside in low-density regions relative to normal points. LOF estimates the local density of a point by comparing it with the densities of its neighbors. It is highly effective in detecting anomalies in datasets with non-uniform density distributions, where global models may fail. However, LOF is computationally expensive complexity $O(n^{2})$ in the worst case and sensitive to the choice of neighborhood size, requiring careful hyperparameter tuning.
A more efficient density estimation approach is HBOS, which models feature distributions independently using histograms. This makes it computationally extremely fast $O(n)$ and scalable to large datasets. However, HBOS assumes feature independence, limiting its effectiveness when strong feature correlations exist. In such cases, its effectiveness diminishes as it fails to capture intricate relationships between features, potentially leading to suboptimal anomaly detection performance.
Isolation-based approaches, such as iForest, take a different perspective by recursively partitioning the feature space. Since anomalies are typically isolated with fewer splits, iForest identifies them based on the depth required to isolate each point. iForest is computationally efficient $O(nlogn)$ and performs well in high-dimensional spaces compared to distance-based methods, but it is struggle with local anomalies. An extension of iForest, INNE, replaces axis-aligned splits with hypersphere-based partitions. This enhances robustness in detecting anomalies in complex distributions, particularly local anomalies.
Statistical approaches model the underlying distribution of data and identify anomalies as points that significantly deviate from expected behavior. ECOD estimates anomaly scores based on the empirical cumulative distribution function (ECDF) for each feature independently. It is parameter-free and computationally efficient $O(n)$, making it highly scalable. However, like HBOS, ECOD assumes feature independence, which can limit its effectiveness in multivariate settings. COPOD improves upon ECOD by leveraging copula functions to model dependencies between features, making it more effective for detecting anomalies in correlated data. However, this comes at the cost of increased computational complexity, making COPOD less scalable for very large datasets.
Table 6: t-SNE visualization of embeddings from 8 models across 6 datasets. Blue points represent normal instances, while red points denote anomalies.
Embeddings Email-Spam SMS-Spam COVID-Fake LIAR2 Hate-Speech OLID
BERT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MINILM [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
O-ada [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
O-small [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
O-large [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Llama [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
stella [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Qwen [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix G Embedding Analysis
To better understand how different embedding models encode normal and anomalous instances, we visualize their embedding spaces using t-SNE projections across 6 datasets. Figure 6 presents the t-SNE plots for embeddings extracted from 8 embedding models, blue points represent normal instances, while red points denote anomalies.
Separation of Normal and Anomalous Instances. As defined in Section A, anomalies should ideally exhibit significant deviation from normal instances in the embedding space. The extent to which embeddings separate anomalies from normal data is a crucial factor in determining their effectiveness for anomaly detection.
Most embedding models exhibits clear separation, particularly in the Email Spam dataset, where anomalous points form distinct regions away from the normal distribution. BERT struggles with clear separation, with many anomalies still embedded within normal clusters. This indicates that these models may not encode sufficient discriminative features for anomaly detection tasks.
Dataset-Specific Challenges. The effectiveness of embedding-based anomaly detection varies significantly across datasets, highlighting the influence of domain characteristics:
  • Spam Detection (Email Spam, SMS Spam): most embedding models perform well, reflecting their ability to capture explicit spam patterns (e.g., domain-specific keywords, unusual syntax). In contrast, BERT shows more overlap between spam and normal messages, leading to weaker anomaly separation.
  • Fake News Detection (COVID-Fake, LIAR2): The separation of anomalies is less pronounced across most embeddings, likely due to the subtle and nuanced nature of misinformation. This suggests that effective detection may require external knowledge or factual reasoning beyond what standard embeddings can provide.
  • Offensive Language (Hate Speech, OLID): All embeddings perform poorly, with anomalies scattered among normal instances. This suggests that hate speech and offensive language often depend on implicit contextual cues rather than explicit linguistic differences, making them harder to distinguish using standard embeddings.
Clustered Anomalies in Spam Detection. For both Email Spam and SMS Spam datasets, the anomalies tend to form compact clusters rather than being scattered as isolated points. This behavior contrasts with other datasets, where anomalies are often more dispersed.
Unlike anomalies in misinformation or hate speech detection, which can manifest in subtle linguistic variations, spam messages tend to exhibit repetitive patterns, including URLs, phone numbers, irregular word spacing and excessive punctuation. Since these patterns are highly distinct but internally consistent, embeddings may cluster them into a well-defined anomaly group rather than spreading them across the feature space.
Appendix H Experiments Environment
The entire pipeline, including embedding extraction and anomaly detection, was implemented in Python 3.9. Experiments were executed on a computational setup equipped with a Ryzen 9 5900X 12-core CPU for data preprocessing and model orchestration, and an Nvidia RTX 3060 GPU with 12GB of memory for model inference and embedding generation.